Sasha has 15 coins worth $2.40. If the coins are all quarters and dimes, how many quarters and dimes does she have?

Answer by math_helper(2461)

d = # dimes
q = # quarters
d+q = 15 (1) (given)
10d + 25q = 240 (2) (given, expressed in cents)

10d + 10q = 150 (eqn(1) x10, call it (1'))
15q = 90 (subtracted (1') from (2))
q = 6 (divided both sides by 15)
Plug in q=6 into (1) ((2) also would work but not as simple of an equation): d=15-6=9

Answer: 9 dimes, 6 quarters

Check: 9+6=15 (ok on # of coins)
10*9 + 25*6 = 90+150 = 240 (ok on value)